The diameters of the outer and inner shells of a semi circular right prism are 30m and 28m respectively. The lenght of the prism is 120m
1.Determine the area of the base in terms of pie

Answer by Alan3354(69443)

The diameters of the outer and inner shells of a semi circular right prism are 30m and 28m respectively. The lenght of the prism is 120m
1.Determine the area of the base in terms of pie
--------------
r = 15 and 14
---
For a semi-circle, area = pi*r^2/2 (Notice it's the Greek letter pi, not the dessert, pie)
---
Area = pi*15^2 - pi*14^2 = 29pi sq meters
That's the area of the semi-circular base.
The length is not a factor.